EA Sports have asked you (an analyst) to find out the mean and standard deviation of the number of minutes people take to complete level 1 of their new game using the below information: The number of minutes people spend playing level 1 follows a normal distribution The probability of player playing for less than 5 minutes is 0.0045 The probability of someone taking less than 15 minutes to complete level 1 is 0.9641


1
Expert's answer
2022-06-02T17:17:43-0400
"P(X<5)=P(Z<\dfrac{5-\mu}{\sigma})=0.0045"

"\dfrac{5-\mu}{\sigma}=-2.612054"


"P(X<15)=P(Z<\dfrac{15-\mu}{\sigma})=0.9641"

"\dfrac{15-\mu}{\sigma}=1.800384"

"\dfrac{5-\mu}{15-\mu}=\dfrac{-2.612054}{1.800384}"

"\mu=\dfrac{5(1.800384)+15(2.612054)}{1.800384+2.612054}"

"\mu=10.92"

"\sigma=\dfrac{15-10.92}{1.800384}"

"\sigma=2.266"